how do you solve the equation n/20=3/260?
Answer by ReadingBoosters(3246)

To solve the equation
n/20 = 3/260
Multiply by sides by 20
n = 3/260(20)
n = 3/13
...............................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.